Fine the equation of the line containing the following two points: (0,-10) and (-5,0)
Answer by lwsshak3(11628)

Fine the equation of the line containing the following two points: (0,-10) and (-5,0)
***
Standard form of an equation for a straight line: y=mx+b, m=slope, b= y-intercept
From given two points:
slope=∆y/∆x=y1-y2/x1-x2=(0-(-10))/(-5-0)=10/-5=-2
Equation:
y=-2x+b
solving for b using one of given points (-5,0)
0=-2*-5+b
0=10+b
b=-10
Equation of line containing points (0,-10) and (5,0)
y=-2x-10